OpenStudy (anonymous):

Question 5 (Multiple Choice Worth 1 points) What is the slope of the line between (-4, 4) and (-1, -2)? 1 2 -2 -1 Question 6 (Multiple Choice Worth 1 points) What is the slope of the line x = -5? undefined 0 -5 5 Question 7 (Multiple Choice Worth 2 points) Select the correct inequality for the graph below. A dotted line passing through points (0, 2) and (1, 0) has shading below. y > -2x + 2 y ≥ -2x + 2 y < -2x + 2 y ≤ -2x + 2

OpenStudy (anonymous):

For the first question is pretty easy. You could directly plot and do a short calculation, since the slope is equal to dy/dx or diviation of y divided by deviation of x. In other word, slope is \[\Delta y/\Delta x\] So that, from (-4, 4) and (-1, -2) you could make a triangle as shown below,  Then determine the deviation on y [ 4-(-2)= 6 ] and deviation on x [-2-(-4)=2 ]. The slope is \[\Delta y/\Delta x=6/2=3\]